In November, 2% more new cars were sold than in October, but 2% fewer new cars were sold in December than in November. If 250,000 new cars were sold in October, how many were sold in December?
(1) 250,000 (2) 249,900 (3) 248,000 (4) 247,750

Answer by Theo(13342)

october sold 250,000 cars

november sold 250,000 + (.02)*250,000 = 250,000 * (1 + .02) = 250,000 * (1.02) = 255,000 cars.

december sold 255,000 - (.02*255,000) = 255,000 * (1 - .02) = 255,000 * (.98) = 249,900 cars.

answer is selection (2).